# Instability of regular planar black holes in four dimensions arising from an infinite sum of curvature corrections

### v1

@@ I Introduction
In the presence of specific matter sources, numerous approaches have been proposed to construct nonsingular BH solutions in four dimensions. One well-known example is nonlinear electrodynamics (NED) within GR. The NED Lagrangian ${\cal L}$ is a nonlinear function of the electromagnetic field strength $F=-F_{\mu\nu}F^{\mu\nu}/4$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the Maxwell tensor associated with a vector field $A_{\mu}$. With suitable choices of the functional form of ${\cal L}(F)$, it is possible to construct SSS charged BH solutions that are free from curvature singularities at their centers [5, 6, 7, 8]. However, such nonsingular BHs are ruled out by angular Laplacian instabilities of vector-field perturbations near their centers [9, 10].
If we consider a canonical scalar field $\phi$ minimally coupled to gravity on the SSS background, the resulting solution is the Schwarzschild BH without scalar hair [11, 12]. This no-hair property holds for a broad class of scalar-tensor theories—including k-essence [13], nonminimally coupled scalar fields with the Ricci scalar $R$ of the form $G_{4}(\phi)R$ [14, 15, 16], and covariant Galileons [17, 18, 19]. A notable exception is the Gauss-Bonnet (GB) curvature invariant $R_{\rm GB}^{2}$ coupled to the scalar field in the form $\xi(\phi)R_{\rm GB}^{2}$ [20, 21, 22, 23, 24, 25, 26, 27, 28]. In this case, the background geometry is modified from the Schwarzschild solution by the presence of a nonvanishing scalar-field profile, but a curvature singularity still remains at the center of BHs.
@@
In four-dimensional spacetime ($d=4$), there have been attempts to extract higher-curvature corrections by employing various regularization schemes [47, 48, 49, 50, 51, 52]. If we restrict ourselves to local gravitational theories involving derivatives of the metric tensor $g_{\mu\nu}$ up to second order in order to avoid Ostrogradsky instabilities, the only nonvanishing Lagrangians in four dimensions are the cosmological constant $\Lambda$ and the Ricci scalar $R$ [53, 54]. The GB term $R_{\rm GB}^{2}$ is a specific combination of quadratic curvature invariants that leads to second-order field equations in arbitrary dimensions, but its contribution to the equations of motion identically vanishes in four dimensions. By rescaling the GB coupling constant $\alpha$ to $\alpha/(d-4)$ and then taking the limit $d\to 4$, one can extract the contribution of the GB term in four dimensions [47]. However, this 4-dimensional Einstein-Gauss-Bonnet (4DEGB) gravity is plagued by several problems—such as the divergences in the perturbation equations [55, 56] and the absence of covariant equations for a graviton [57, 58].
@@
In theories arising from an infinite sum of curvature corrections, the geometry of planar BHs is modified by the presence of a nonvanishing scalar-field radial derivative of the form $\phi^{\prime}(r)=1/r$ [52]. The metric components of such BHs remain regular at $r=0$, but it is not yet clear whether they are stable against linear perturbations. In this paper, we address the linear stability of planar BHs that are regular at $r=0$. Since the two-dimensional topology of these BHs differs from that of the SSS spacetime, it is necessary to formulate BH perturbations on a different background within the framework of shift-symmetric Horndeski theories. We compute the second-order actions for both the odd- and even-parity sectors and derive the linear stability conditions for the three dynamical degrees of freedom: one scalar mode and two gravitational modes. In the odd-parity sector, we show that the BH solutions with $\phi^{\prime}(r)=1/r$ suffer from ghost and Laplacian instabilities near $r=0$. Moreover, the kinetic term of the even-parity scalar-field perturbation vanishes both inside and outside the horizon. As a result, BHs that are regular at $r=0$ are plagued by an infinitely strong coupling problem. Thus, these regular planar BHs are ruled out as stable configurations.
@@ II Regularized four-dimensional Lovelock gravity
$G_{2}(X)$ $=$ $-\Lambda+\frac{4\ell^{2}X^{2}(1+6\ell^{2}X)}{(1-2\ell^{2}X)^{3}},$
$G_{3}(X)$ $=$ $\frac{1-10\ell^{2}X}{(1-2\ell^{2}X)^{3}}\,,$
$G_{4}(X)$ $=$ $\frac{1}{2(1-2\ell^{2}X)^{2}}\,,$
$G_{5}(X)$ $=$ $-2\ell^{2}\left[\frac{1-2\ell^{4}X^{2}}{(1-2\ell^{2}X)^{2}}+\ln\left(\frac{2\ell^{2}|X|}{1-2\ell^{2}X}\right)\right].$ (13)
It should be noted that this action is well-defined only when $01-2\ell^{2}X>$ 0. If this condition is not satisfied, the last term of $G_{5}(X)$ can be generalized to the form $\ln[2\ell^{2}|X|/|1-2\ell^{2}X|]$. In particular, $X$ is not generally bounded in this theory. A discussion of this generalization of the Horndeski coupling functions is presented in Appendix A.
@@ III.1 Model 1
From this expression we can read off an effective cosmological constant $\Lambda_{\rm eff}=-3/(\ell_{\Lambda}^{2}-\ell^{2})$, and an ADM mass $\mu=M\ell_{\Lambda}^{4}/(\ell_{\Lambda}^{2}-\ell^{2})^{2}$. For the BH to be timelike at spatial infinity, we require the condition $\ell_{\Lambda}^{2}>\ell^{2}>0$, or, equivalently, for $-3/\ell^{2}<\Lambda<0$, the variable $r$ is spacelike. [^2] For the fine-tuned case $\Lambda=-3/\ell^{2}$, the metric component is given by $h=r^{5}[c_{1}-1/(\ell^{2}r^{3})]$, where $c_{1}$ is a positive constant.
In the limits $r\to 0$ and $r\to\infty$, the scalar-field kinetic term $X=-(1/2)h\phi^{\prime 2}$ approaches $1/(2\ell^{2})$ and $-1/[2(\ell_{\Lambda}^{2}-\ell^{2})]$, respectively. As long as $X$ is a continuously decreasing function of $r$, it should be in the range
$-\frac{1}{2(\ell_{\Lambda}^{2}-\ell^{2})}<X<\frac{1}{2\ell^{2}}\,.$ (32)
Note that $X$ crosses 0 at the horizon ($h=0$). In Fig. 1, we plot $\ell^{2}X$ and $f\,(=h)$ versus $r/\ell$ in the range $r\geq 0$ for $M=\ell_{\Lambda}$ and $\ell=\ell_{\Lambda}/2$. The quantity $X$ decreases smoothly from $1/(2\ell^{2})$ at $r=0$ to $-1/[2(\ell_{\Lambda}^{2}-\ell^{2})]$ at spatial infinity. For $r<r_{h}$, the metric function is negative ($f<0$), whereas it becomes positive ($f>0$) for $r>r_{h}$. At $r=r_{h}$, both $X$ and $f$ vanish.
Figure 1: The metric function $f~{}(=h)$ and the rescaled scalar derivative $\ell^{2}X$ are plotted as functions of $r/\ell$ in the range $r\geq 0$ for Model 1 with $M=\ell_{\Lambda}$ and $\ell=\ell_{\Lambda}/2$. The horizon is located at $r_{h}=2^{1/3}\ell_{\Lambda}=2^{4/3}\ell$, where both $f$ and $X$ vanish. The kinetic term decreases smoothly from the value $X=1/(2\ell^{2})$ at $r=0$ to the asymptotic value $X=-1/[2(\ell_{\Lambda}^{2}-\ell^{2})]$, as $r\to\infty$.
@@ III.2 Model 2
which approaches the finite value $X\to 1/(2\ell^{2})$ for $r\to+0$. In the other limit $r\to\infty$, $X$ approaches $-(1/2\ell^{2})\tanh(\ell^{2}/\ell_{\Lambda}^{2})$. Therefore, $X$ is bounded in the range
$-\frac{1}{2\ell^{2}}\tanh\left(\frac{\ell^{2}}{\ell_{\Lambda}^{2}}\right)<X<\frac{1}{2\ell^{2}}\,.$ (43)
@@ III.3 Models corresponding to each value of $n$
×$\times(r\phi^{\prime}-1)^{2}(f^{\prime}-2f\phi^{\prime})\,,$
@@ IV Perturbations around the planar BH
We will use this form of $Z$ together with the normalization condition (48) for deriving the second-order action of perturbations. We have restricted the range of $x_{2}$ to $-\pi\leq x_{2}\leq\pi$, as in the cases of cylindrical and toroidal BHs. However, the local stability conditions remain the same as those for planar BHs. We will further consider a finite range of $x_{3}$, for instance, $0\leq x_{3}\leq 1$, without loss of generality. We note that the mode function here is simpler than the spherical harmonics $Y_{lm}(\theta,\varphi)$ used in the analysis for BH perturbations on the SSS background [74, 75].
@@ V Odd-parity perturbations
~$\tilde{{\cal L}}_{\rm odd}$ $=$ $\frac{1}{4}\sqrt{\frac{h}{f}}{\cal H}\left[2\chi\left(\dot{W}-Q^{\prime}+\frac{2Q}{r}\right)-\chi^{2}\right]$ (64)
@@
Varying ~$\tilde{{\cal L}}_{\rm odd}$ with respect to the Lagrange multiplier $\chi$ yields
@@
Substituting Eq. (65) into Eq. (64), we find that ~$\tilde{{\cal L}}^{\rm odd}$ is equivalent to ${\cal L}^{\rm odd}$. Varying Eq. (64) with respect to $W$ and $Q$, we obtain
@@ V.2 Application to concrete models
Let us examine the linear stability of regular planar BHs discussed in Sec. III. In all these models, we have shown the existence of a background solution characterized by $\phi^{\prime}(r)=1/r$. For this solution, the metric functions satisfy $f=h$. Using these properties, the difference between ${\cal F}$ and ${\cal G}$ is given by
@@ V.2.1 Model 1
×$\qquad\qquad\,\times[2M\ell^{2}\ell_{\Lambda}^{2}+(\ell_{\Lambda}^{2}-\ell^{2})r^{3}]/(\ell_{\Lambda}^{4}r^{6})\,,$ (85)
@@
Since the leading-order terms of ${\cal F}$ and $c_{\Omega,{\rm odd}}^{2}$ are negative, both ghost and Laplacian instabilities arise near $r=0$. This shows that the BH solution (29) with $\phi^{\prime}(r)=1/r$ is unstable due to the rapid growth of odd-parity perturbations around $r=0$.
@@ V.2.2 Model 2
×$\qquad\qquad\times\left[12M\ell^{2}\sinh(x)-r^{3}\cosh(x)\right]\,,$ (89)
@@
which are both negative. Therefore, the BH solution (40) is unstable due to the presence of both ghost and Laplacian instabilities near $r=0$. The difference from Model 1 is that $c_{\Omega,{\rm odd}}^{2}$ diverges to $\infty$- ∞ in the limit $r\to+0$, indicating that the Laplacian instability is even more severe in Model 2.
In both Models 1 and 2, we have shown that planar BHs regular at $r=0$ are ruled out due to the instability of odd-parity perturbations. In contrast, in GR, we have $G_{4}=1/2$ and $G_{5}=0$, which leads to ${\cal F}={\cal G}=1$ and $c_{\Omega,{\rm odd}}^{2}=1$. In models involving an infinite sum over $n$, the function $G_{4}(X)$ exhibits a nontrivial dependence on $X$, as seen in Eqs. (13) and (15). This specific feature leads to negative values of ${\cal F}$ and $c_{\Omega,{\rm odd}}^{2}$ near $r=0$.
@@ VI Even-parity perturbations
where $a_{4}$ is related to ${\cal H}$ by $a_{4}=(f/2)\sqrt{h/f}\,{\cal H}$. In the following, we derive the no-ghost conditions and radial propagation speeds for the dynamical perturbations $\psi$ and $\delta\phi$. We then apply these results to analyze the stability of BHs discussed in Sec. III. Since the no-ghost conditions and radial speeds are sufficient to rule out all these BH solutions, we do not derive the angular propagation speeds in the even-parity sector. In GR without a scalar field $\phi$, we discuss the linear stability of planar BHs against perturbations in both radial and angular directions in Appendix C. In this case, all the stability conditions are trivially satisfied.
@@ VI.1 Region with $f>0$, $h>0$
$K_{11}$ $=$ $\frac{{\cal P}\mu^{4}f^{2}h^{2}}{2\sqrt{fh}\,{\cal H}^{2}[2rfk^{2}{\cal H}+h\mu(rf^{\prime}-2f)]^{2}}>0\,,$ (102)
${\rm det}\,{\bm{K}}$ $=$ $K_{11}K_{22}-K_{12}^{2}$ (103)
@@
$c_{r1,{\rm even}}^{2}$ $=$ $\frac{{\cal G}}{{\cal F}}\,,$ (106)
$c_{r2,{\rm even}}^{2}$ $=$ $\frac{1}{(2{\cal P}-{\cal F})\mu^{2}\sqrt{fh}}\left[\sqrt{\frac{h}{f}}(\mu-2r{\cal H})(r^{2}f^{\prime}{\cal H}^{2}+2rf{\cal G}{\cal H}-f\mu{\cal G})+4c_{2}r^{2}\phi^{\prime}{\cal H}^{2}+8c_{3}h\phi^{\prime}r^{2}{\cal H}(\mu-r{\cal H})\right]\,.$ (107)
@@ VI.2 Region with $f<0$, $h<0$
$c_{r1,{\rm even}}^{2}$ $=$ $\frac{{\cal F}}{{\cal G}}\,,$ (111)
$c_{r2,{\rm even}}^{2}$ $=$ $\frac{fh\mu^{2}(2{\cal P}-{\cal F})}{h(\mu-2r{\cal H})(r^{2}f^{\prime}{\cal H}^{2}+2rf{\cal G}{\cal H}-f\mu{\cal G})-4{\cal H}[{\cal H}c_{2}-2c_{3}h(r{\cal H}-\mu)]r^{2}\phi^{\prime}\sqrt{fh}}\,.$ (112)
Both $c_{r1,{\rm even}}^{2}$ and $c_{r2,{\rm even}}^{2}$ must to be positive to avoid Laplacian instabilities in the radial direction. Notice that $c_{r1,{\rm even}}^{2}$ corresponds to the squared radial propagation speed of the gravitational perturbation $\psi$. Indeed, this is equivalent to the value $c_{r,{\rm odd}}^{2}$ in the region $f<0$ and $h<0$. As long as the linear stability conditions ${\cal F}>0$ and ${\cal G}>0$ in the odd-parity sector are satisfied, the Laplacian instability for $\psi$ is avoided. The scalar-field perturbation $\delta\phi$ has the squared radial propagation speed $c_{r2,{\rm even}}^{2}$ given by Eq. (112). Compared to the expression for $c_{r2,{\rm even}}^{2}$ in the region $f>0$ and $h>0$, where the term $2{\cal P}-{\cal F}$ appears in the denominator of Eq. (107), the same term appears in the numerator of Eq. (112).

### v2

@@ I Introduction
[…]
If we consider a canonical scalar field $\phi$ minimally coupled to gravity on the SSS background, the resulting solution is the Schwarzschild BH without scalar hair [11, 12]. This no-hair property holds for a broad class of scalar-tensor theories—including k-essence [13], nonminimally coupled scalar fields with the Ricci scalar $R$ of the form $G_{4}(\phi)R$ [14, 15, 16], and covariant Galileons [17, 18, 19]. A notable exception is the Gauss-Bonnet (GB) curvature invariant $R_{\rm GB}^{2}$ coupled to the scalar field in the form $\xi(\phi)R_{\rm GB}^{2}$ [20, 21, 22, 23, 24, 25, 26, 27, 19, 28]. In this case, the background geometry is modified from the Schwarzschild solution by the presence of a nonvanishing scalar-field profile, but a curvature singularity still remains at the center of BHs.
@@
In four-dimensional spacetime ($d=4$), there have been attempts to extract higher-curvature corrections by employing various regularization schemes [47, 48, 49, 50, 51, 52] (see also Ref. [53]). If we restrict ourselves to local gravitational theories involving derivatives of the metric tensor $g_{\mu\nu}$ up to second order in order to avoid Ostrogradsky instabilities, the only nonvanishing Lagrangians in four dimensions are the cosmological constant $\Lambda$ and the Ricci scalar $R$ [54, 55]. The GB term $R_{\rm GB}^{2}$ is a specific combination of quadratic curvature invariants that leads to second-order field equations in arbitrary dimensions, but its contribution to the equations of motion identically vanishes in four dimensions. By rescaling the GB coupling constant $\alpha$ to $\alpha/(d-4)$ and then taking the limit $d\to 4$, one can extract the contribution of the GB term in four dimensions [47]. However, this 4-dimensional Einstein-Gauss-Bonnet (4DEGB) gravity is plagued by several problems—such as the divergences in the perturbation equations [56, 57] and the absence of covariant equations for a graviton [58, 59].
@@
In theories arising from an infinite sum of curvature corrections, the geometry of planar BHs is modified by the presence of a nonvanishing scalar-field radial derivative of the form $\phi^{\prime}(r)=1/r$ [52]. We emphasize that this type of solution arises in planar two-dimensional geometry, whereas for standard spherically symmetric BHs with a two-dimensional closed geometry, the scalar-field solution takes a different form. In the latter case, no linearly stable, regular, spherically symmetric BHs with nontrivial scalar hair have been reported in the literature [19, 28].
The metric components of planar BHs found in Ref. [52] remain regular at $r=0$, but it is not yet clear whether they are stable against linear perturbations. In this paper, we address the linear stability of planar BHs that are regular at $r=0$. Since the two-dimensional topology of these BHs differs from that of the SSS spacetime, it is necessary to formulate BH perturbations on a different background within the framework of shift-symmetric Horndeski theories. We compute the second-order actions for both the odd- and even-parity sectors and derive the linear stability conditions for the three dynamical degrees of freedom: one scalar mode and two gravitational modes. In the odd-parity sector, we show that the BH solutions with $\phi^{\prime}(r)=1/r$ suffer from ghost and Laplacian instabilities near $r=0$. Moreover, the kinetic term of the even-parity scalar-field perturbation vanishes both inside and outside the horizon. As a result, BHs that are regular at $r=0$ are plagued by an infinitely strong coupling problem. Thus, these regular planar BHs are ruled out as stable configurations.
@@ II Regularized four-dimensional Lovelock gravity
$\hskip-28.45274ptG_{2}(X)$ $=$ $-\Lambda+\frac{4\ell^{2}X^{2}(1+6\ell^{2}X)}{(1-2\ell^{2}X)^{3}},$
$\hskip-28.45274ptG_{3}(X)$ $=$ $\frac{1-10\ell^{2}X}{(1-2\ell^{2}X)^{3}}\,$,
$\hskip-28.45274ptG_{4}(X)$ $=$ $\frac{1}{2(1-2\ell^{2}X)^{2}}\,$,
$\hskip-28.45274ptG_{5}(X)$ $=$ $-2\ell^{2}\left[\frac{1-2\ell^{4}X^{2}}{(1-2\ell^{2}X)^{2}}+\ln\left(\frac{2\ell^{2}|X|}{1-2\ell^{2}X}\right)\right].$ (13)
It should be noted that this action is well-defined only when $1-2\ell^{2}X>0$. If this condition is not satisfied, the last term of $G_{5}(X)$ can be generalized to the form $\ln[2\ell^{2}|X|/|1-2\ell^{2}X|]$. In particular, $X$ is not generally bounded in this theory. A discussion of this generalization of the Horndeski coupling functions is presented in Appendix A.
@@ III.1 Model 1
From this expression we can read off an effective cosmological constant $\Lambda_{\rm eff}=-3/(\ell_{\Lambda}^{2}-\ell^{2})$, and an ADM mass $\mu=M\ell_{\Lambda}^{4}/(\ell_{\Lambda}^{2}-\ell^{2})^{2}$. For the BH to be timelike at spatial infinity, we require the condition $\ell_{\Lambda}^{2}>\ell^{2}>0$, or, equivalently, for $3/\ell^{2}<\Lambda<0$, the variable $r$ is spacelike. [^2] For the fine-tuned case $\Lambda=-3/\ell^{2}$, the metric component is given by $h=r^{5}[c_{1}-1/(\ell^{2}r^{3})]$, where $c_{1}$ is a positive constant.
In the limits $r\to 0$ and $r\to\infty$, the scalar-field kinetic term $X=-(1/2)h\phi^{\prime 2}$ approaches $1/(2\ell^{2})$ and $1/[2(\ell_{\Lambda}^{2}-\ell^{2})]$, respectively. As long as $X$ is a continuously decreasing function of $r$, it should be in the range
$\frac{1}{2(\ell_{\Lambda}^{2}-\ell^{2})}<X<\frac{1}{2\ell^{2}}\,.$ (32)
Note that $X$ crosses 0 at the horizon ($h=0$). In Fig. 1, we plot $\ell^{2}X$ and f(=h)f\,(=h) versus $r/\ell$ in the range $r\geq 0$ for $M=\ell_{\Lambda}$ and $\ell=\ell_{\Lambda}/2$. The quantity $X$ decreases smoothly from $1/(2\ell^{2})$ at $r=0$ to $1/[2(\ell_{\Lambda}^{2}-\ell^{2})]$ at spatial infinity. For $r<r_{h}$, the metric function is negative ($f<0$), whereas it becomes positive ($f>0$) for $r>r_{h}$. At $r=r_{h}$, both $X$ and $f$ vanish.
Figure 1: The metric function f(=h)f~(=h) and the rescaled scalar derivative $\ell^{2}X$ are plotted as functions of $r/\ell$ in the range $r\geq 0$ for Model 1 with $M=\ell_{\Lambda}$ and $\ell=\ell_{\Lambda}/2$. The horizon is located at $r_{h}=2^{1/3}\ell_{\Lambda}=2^{4/3}\ell$, where both $f$ and $X$ vanish. The kinetic term decreases smoothly from the value $X=1/(2\ell^{2})$ at $r=0$ to the asymptotic value $X=-1/[2(\ell_{\Lambda}^{2}-\ell^{2})]$, as $r\to\infty$.
@@ III.2 Model 2
which approaches the finite value $X\to 1/(2\ell^{2})$ for $r\to+0$. In the other limit $r\to\infty$, $X$ approaches $(1/2\ell^{2})\tanh(\ell^{2}/\ell_{\Lambda}^{2})$. Therefore, $X$ is bounded in the range
$\frac{1}{2\ell^{2}}\tanh\left(\frac{\ell^{2}}{\ell_{\Lambda}^{2}}\right)<X<\frac{1}{2\ell^{2}}\,.$ (43)
@@ III.3 Models corresponding to each value of $n$
$\times(r\phi^{\prime}-1)^{2}(f^{\prime}-2f\phi^{\prime})\,$,
@@ IV Perturbations around the planar BH
We will use this form of $Z$ together with the normalization condition (48) for deriving the second-order action of perturbations. We have restricted the range of $x_{2}$ to $\pi\leq x_{2}\leq\pi$, as in the cases of cylindrical and toroidal BHs. However, the local stability conditions remain the same as those for planar BHs. We will further consider a finite range of $x_{3}$, for instance, $0\leq x_{3}\leq 1$, without loss of generality. We note that the mode function here is simpler than the spherical harmonics $Y_{lm}(\theta,\varphi)$ used in the analysis for BH perturbations on the SSS background [75, 76].
@@ V Odd-parity perturbations
$\tilde{{\cal L}}_{\rm odd}$ $=$ $\frac{1}{4}\sqrt{\frac{h}{f}}{\cal H}\left[2\chi\left(\dot{W}-Q^{\prime}+\frac{2Q}{r}\right)-\chi^{2}\right]$ (64)
@@
Varying $\tilde{{\cal L}}_{\rm odd}$ with respect to the Lagrange multiplier $\chi$ yields
@@
Substituting Eq. (65) into Eq. (64), we find that $\tilde{{\cal L}}^{\rm odd}$ is equivalent to ${\cal L}^{\rm odd}$. Varying Eq. (64) with respect to $W$ and $Q$, we obtain
@@ V.2 Application to concrete models
[…]
@@ V.2.1 Model 1
$\qquad\qquad\,\times[2M\ell^{2}\ell_{\Lambda}^{2}+(\ell_{\Lambda}^{2}-\ell^{2})r^{3}]/(\ell_{\Lambda}^{4}r^{6})\,$, (85)
@@
[…]
@@ V.2.2 Model 2
$\qquad\qquad\times\left[12M\ell^{2}\sinh(x)-r^{3}\cosh(x)\right]\,$, (89)
@@
which are both negative. Therefore, the BH solution (40) is unstable due to the presence of both ghost and Laplacian instabilities near $r=0$. The difference from Model 1 is that $c_{\Omega,{\rm odd}}^{2}$ diverges to -$\infty$ in the limit $r\to+0$, indicating that the Laplacian instability is even more severe in Model 2.
[…]
@@ VI Even-parity perturbations
[…]
@@ VI.1 Region with $f>0$, $h>0$
$\hskip-22.76228ptK_{11}$ $=$ $\frac{{\cal P}\mu^{4}f^{2}h^{2}}{2\sqrt{fh}\,{\cal H}^{2}[2rfk^{2}{\cal H}+h\mu(rf^{\prime}-2f)]^{2}}>0\,$, (102)
$\hskip-22.76228pt{\rm det}\,{\bm{K}}$ $=$ $K_{11}K_{22}-K_{12}^{2}$ (103)
@@
$\hskip-14.22636ptc_{r1,{\rm even}}^{2}$ $=$ $\frac{{\cal G}}{{\cal F}}\,$, (106)
$\hskip-14.22636ptc_{r2,{\rm even}}^{2}$ $=$ $\frac{1}{(2{\cal P}-{\cal F})\mu^{2}\sqrt{fh}}\left[\sqrt{\frac{h}{f}}(\mu-2r{\cal H})(r^{2}f^{\prime}{\cal H}^{2}+2rf{\cal G}{\cal H}-f\mu{\cal G})+4c_{2}r^{2}\phi^{\prime}{\cal H}^{2}+8c_{3}h\phi^{\prime}r^{2}{\cal H}(\mu-r{\cal H})\right]\,.$ (107)
@@ VI.2 Region with $f<0$, $h<0$
$\hskip-14.22636ptc_{r1,{\rm even}}^{2}$ $=$ $\frac{{\cal F}}{{\cal G}}\,$, (111)
$\hskip-14.22636ptc_{r2,{\rm even}}^{2}$ $=$ $\frac{fh\mu^{2}(2{\cal P}-{\cal F})}{h(\mu-2r{\cal H})(r^{2}f^{\prime}{\cal H}^{2}+2rf{\cal G}{\cal H}-f\mu{\cal G})-4{\cal H}[{\cal H}c_{2}-2c_{3}h(r{\cal H}-\mu)]r^{2}\phi^{\prime}\sqrt{fh}}\,.$ (112)
[…]